MATH 367, spg 98 PROJECT 2 Due: Mon Mar 2

NAME:

PROBLEM: You are in a game show, where you can win $10,000. The game is as follows: there are 3 doors, which are closed. Behind only one of them there is a box containing the $10,000 in $100 bills. All you have to do is to choose one of these doors. Once you made your choice, the host of the show puts you in a quandary by opening one of the other two reamining doors. (Since he knows where the money is, he opens a door without the money behind it). Now you see that there are two remaining closed doors. And the host gives you the opportunity of changing your choice to the other unopened door.

1. What should you do? Should you change doors or not? Another way of phrasing this question: what is the best winning strategy (a) to switch or (b) not to switch?

2. Justify your decision in 1. Give me a proof that your answer is the best strategy.